# Emulation of Complex Matrix Multiplication based on the Chinese Remainder Theorem

### v1

@@ III-B Scaling Vector Determination
First, we convert $\hat{A}$ and $\hat{B}$ to integer matrices $\hat{A}^{\prime}$ and $\hat{B}^{\prime}$ as
@@
We also consider two strategies: a fast mode and an accurate mode. From $|\hat{a}^{\prime}_{ij}|\leq|\hat{a}_{ij}|\cdot\mu_{i}$ and $|\hat{b}^{\prime}_{ij}|\leq|\hat{b}_{ij}|\cdot\nu_{j}$,
@@
From (6), the row norms of $\hat{A}$ satisfy $\|\hat{a}_{i,:}\|_{2}=\|\hat{a}_{i+m,:}\|_{2}$ for $1\leq i\leq m$, and the column norms of $\hat{B}$ satisfy $\|\hat{b}_{:,j}\|_{2}=\|\hat{b}_{:,j+n}\|_{2}$ for $1\leq j\leq n$. Thus, the scaling values of rows $i$ and $i+m$ of $\hat{A}$ are identical, and the same holds for columns $j$ and $j+n$ of $\hat{B}$. Consequently, the scaling vectors reduce to $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$. The elements of $\mu$ and $\nu$ can then be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2], i.e.,
@@
For any real number $x\in\mathbb{R}$, $\lfloor\log_{2}|x|\rfloor$ represents the Unit in the First Place (UFP) of $x$, and thus it can be obtained directly using the $\mathtt{ilogb}(\cdot)$ function. Following [uchino_ozaki2], the quantity $(\log_{2}(P-1)-1)/2-1$ is precomputed as $\mathcal{P}^{\prime}_{\mathit{fast}}:=\mathrm{single}(\log_{2}(P-1)/2-1.5)$. Furthermore, $\log_{2}\|\hat{a}_{i,:}\|_{2}$ and $\log_{2}\|\hat{b}_{:,j}\|_{2}$ can be upper-bounded by $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{a}_{ih}^{2})$ and $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{b}_{hj}^{2})$, respectively, using floating-point arithmetic in round-up mode. Here, $\mathtt{\_\_log2f}$ denotes the fast built-in base-2 logarithm function provided by both the CUDA Math API and the HIP math API. According to NVIDIA’s official CUDA documentation, the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is $2^{-22}$ for $x\in[0.5,2]$, and at most 2 Units in the Last Place (ULP) otherwise [cuda]. The definition of ULP used here follows Muller [muller:inria-00070503]. In addition, the AMD HIP documentation states that the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is 1 ULP for $x\in[10^{-6},10^{6}]$ [hip], although the precise definition of ULP is not specified. Based on these specifications, we assume the bound $\mathtt{\_\_log2f}(x)\leq(\log_{2}x)/(1-4u)$, where $u$ is the unit roundoff of FP32 arithmetic. If $x$ is not sufficiently close to $1$, it is first scaled to $x^{\prime}:=x/2^{\lfloor\log_{2}|x|\rfloor}$, after which the logarithm is evaluated as $\mathtt{\_\_log2f}(x^{\prime})+\lfloor\log_{2}|x|\rfloor$. Based on the above considerations, for $\delta:=\mathrm{single}(0.5/(1-4u))$ computed in round-down mode, the scaling vectors $\mu$ and $\nu$ are determined as follows:
@@
where $\mathcal{P}^{\prime}_{\mathit{fast}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using fused multiply-add (fma) operation in round-down mode.
In accurate mode, first, we apply diagonal scaling to $A_{R}$, $A_{I}$, $B_{R}$, and $B_{I}$ so that the integer parts of the maximum absolute values in each row of $A_{R}$ and $A_{I}$, and in each column of $B_{R}$ and $B_{I}$, fit within 6 bits. Specifically, we scale them as $\mathrm{diag}(\bar{\mu})A_{R}$, $\mathrm{diag}(\bar{\mu})A_{I}$, $B_{R}\mathrm{diag}(\bar{\nu})$, $B_{I}\mathrm{diag}(\bar{\nu})$, and then round them upward using the ceiling operation to obtain 7-bit or smaller integer matrices $\bar{A}_{R},\ \bar{A}_{I},\ \bar{B}_{R},\ \bar{B}_{I}$, respectively. These matrices satisfy
@@
where no computing error occurs in $\bar{A}_{R}-\bar{A}_{I}$ and $\bar{B}_{R}-\bar{B}_{I}$ in INT8. Then, the elements of $\mu$ and $\nu$ can be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2], i.e.,
@@
where $\mathcal{P}^{\prime}_{\mathit{accu}}:=\mathrm{single}(\log_{2}(P-1)/2-0.5)$, and $\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using fma in round-down mode.
@@ III-C Performance Analysis
To analyze the performance characteristics of the proposed method, we employ a simple yet expressive performance model that captures both computation and memory-transfer costs. Let $b$ denote the peak sustained memory bandwidth (in Byte/s) and $p$ denote the peak throughput of the INT8 matrix-multiplication engine (in OPS). For a given operation, the memory-bound execution time is estimated as the required data movement divided by $b$, while the compute-bound time is obtained as the arithmetic workload divided by $p$. Because even memory-bound kernels incur a non-negligible amount of arithmetic overhead, we include a correction term $c$.
@@ III-C1 Fast Mode
Let the input matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$ be double precision. In step 1, both $A$ and $B$ are loaded twice: once for computing the rowwise maximum absolute values of $\hat{A}$ and the columnwise maximum absolute values of $\hat{B}$, and once for computing the residues of $\hat{A}$ and $\hat{B}$. Afterward, the scaling vectors and the INT8 matrices are stored. Each scaling factor is a power of two, so only its exponent needs to be stored as an INT16. Including the correction term $c$, which accounts for arithmetic overhead even in memory-bound operations, the total execution time of step 1 can be estimated as $((32+3N+c)k+2)(m+n)/b$. In step 2, the INT8 GEMM kernels are executed $3N$ times. The resulting $3N$ INT32 matrices are then loaded and converted into $N$ INT8 complex matrices. Accounting for computational overhead $c$, the execution time of step 2 is estimated as $6Nmnk/p+(14N+c)mn/b$. In step 3, the $N$ INT8 complex matrices are loaded, and the CRT accumulation and final reduction are performed. The scaling vectors are then loaded, the inverse scaling is applied, and finally the resulting double-precision complex matrix is stored. With the correction term $c$, the execution time of step 3 is estimated as $(2(m+n)+(2N+16+c)mn)/b$.
Combining the costs derived for steps 1, 2, and 3, the total execution time of the proposed method can be estimated as the sum of its memory-bound and compute-bound components. Aggregating all terms yields the following expression for the overall execution time:
@@ III-C2 Accurate Mode
and for the single-precision case,
@@
Figs. 2 and 3 show performance-model heatmaps for the fast and accurate modes in single and double precision for $6$ and $13$ moduli, respectively. The TFLOPS was computed as
@@
where $t$ is the execution time in seconds. For example, assuming an effective memory bandwidth of 2–4 TByte/s and an observed INT8 GEMM kernel throughput of 1500 TFLOPS on the GH200, the model predicts that ZGEMM emulation in the accurate mode would achieve approximately 120 TFLOPS.
Figure 2: Performance-model heatmaps for the single-precision complex matrix multiplication emulation. The left panel corresponds to the fast mode and the right panel to the accurate mode. The horizontal axis denotes the attainable memory bandwidth, and the vertical axis denotes the achievable INT8 GEMM throughput. The problem size is fixed to $m=n=k=16384$, and the correction term is set to $c=6$, equal to the number of moduli used in the emulation. The color scale indicates the predicted throughput in TFLOPS of the proposed emulation.
Figure 3: Performance-model heatmaps for the double-precision complex matrix multiplication emulation. The left panel corresponds to the fast mode and the right panel to the accurate mode. The horizontal axis denotes the attainable memory bandwidth, and the vertical axis denotes the achievable INT8 GEMM throughput. The problem size is fixed to $m=n=k=16384$, and the correction term is set to $c=13$, equal to the number of moduli used in the emulation. The color scale indicates the predicted throughput in TFLOPS of the proposed emulation.
@@ III-D GEMM Emulation Library
In this study, we developed a library that provides a complete emulation environment for GEMM operations based on the Ozaki-II scheme. Details of the library and usage instructions are provided in [GEMMul8]. The library supports both CUDA and HIP backends, and the CUDA-based source code is automatically translated into HIP C++ when compiled under the HIP backend. The implementation provides emulation routines for SGEMM and DGEMM as proposed in previous work, as well as the CGEMM and ZGEMM emulation kernels developed in the present study. The supported data types are float, double, cuComplex, cuDoubleComplex, hipFloatComplex, hipDoubleComplex.
To enable transparent integration with existing HPC applications, the library can be activated by adding the path of the shared library to the environment variable LD_PRELOAD. Once activated, the following GEMM calls to CUDA or HIP backends are dynamically redirected to the emulation routines, allowing users to evaluate GEMM emulation without modifying their applications:
@@
The implementation consists of INT8 matrix multiplications, fixed-order basic arithmetic operations, and the evaluation of the $\log_{2}(\cdot)$ function appearing in (11), (12), (13), and (14). Rounding errors arise only in the computation of the scaling vectors defined by (11), (12), (13), and (14), and in the accumulation and final reduction stages of the CRT. Except for the logarithm evaluation, all computations are deterministic sequences of basic arithmetic, and thus the emulation produces bitwise-reproducible results as long as the precision of the $\log_{2}(\cdot)$ computation is identical. When the API versions coincide, each backend employs identical implementations of __log2f, ensuring reproducibility across runs.
@@ IV Numerical Results
Here, we compare the following methods with each other:
@@
The hipBLAS does not provide hipblasCgemm3m or hipblasZgemm3m and the same limitation applies to the rocBLAS, though it is not used in this study. We have empirically confirmed that both the performance and numerical accuracy of cublasCgemm3m are essentially identical to those of cublasCgemm; thus, we omit its results from the following comparisons. The cuBLAS implementation of Ozaki-I-based emulation provides two execution modes: Dynamic Mantissa Control and Fixed Mantissa Control. In the Dynamic Mantissa Control mode, the routine determines the number of mantissa bits, equivalently, the number of slices, required for fixed-point emulation such that the numerical accuracy matches or exceeds that of FP64. If the required mantissa length exceeds either an internal limit or a user-specified maximum, the computation is automatically executed in native FP64. In the Fixed Mantissa Control mode, the user specifies the mantissa-bit length used for the fixed-point representation. This mode reduces runtime overhead by eliminating dynamic precision selection, but it does not ensure FP64-level accuracy. In this study, the Fixed Mantissa Control mode is used to enable a direct comparison of computational throughput and to quantify the performance gain attributable to fixed-precision emulation.
@@ IV-A Accuracy
where $\mathrm{rand}\in(0,1]$ denotes a uniform random number and $\mathrm{randn}$ denotes a standard normal random number. Both types of random numbers were generated using either the cuRAND API or the hipRAND library. Following the methodology adopted in prior studies about the Ozaki schemes [ozaki2012error, ozaki2013generalization, mukunoki2020, ootomo2024dgemm, uchino2025Performance, ozaki-scheme2, uchino_ozaki2], we introduce the parameter $\phi$ to control the dynamic range of the input matrices. In accordance with these studies, we specify $\phi\in\{0,0.5,1,1.5\}$ for single-precision tests and $\phi\in\{0.5,1,2,4\}$ for double-precision tests. As $\phi$ increases, the maximum magnitude grows exponentially, whereas the median and interquartile range remain comparatively stable. This confirms that $\phi$ effectively widens the dynamic range without substantially affecting the central tendency of the distribution. Such controlled variation is essential for evaluating the performance and numerical stability of the Ozaki schemes.
Figs. 4 and 5 show the accuracy of CGEMM and ZGEMM, respectively, observed on the GH200. Similar results were obtained on other GPUs. The maximum relative error in the figures were calculated as
@@
In CGEMM emulation, fast-$N$ with $6\leq N\leq 9$ and accu-$N$ with $6\leq N\leq 8$ provided the accuracy comparable with that of CGEMM. In ZGEMM emulation, fast-$N$ with $13\leq N\leq 18$ and accu-$N$ with $13\leq N\leq 17$ achieved the accuracy comparable with that of ZGEMM. As $\phi$ increases, the range of exponents for $A$ and $B$ expands, making it easier to overestimate the upper bound $2\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ in the condition (4). DGEMM emulation in [uchino_ozaki2] required at least 14 moduli to achieve DGEMM-level accuracy, whereas the proposed method attains ZGEMM-level accuracy with only 13 moduli. This reduction is attributed to the use of the Karatsuba algorithm in the INT8 complex matrix multiplication, which suppresses precision loss more effectively than floating-point arithmetic.
@@
Figure 4: Maximum relative error of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=n=1024$ and $k=16384$.
Figure 5: Maximum relative error of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=n=1024$ and $k=16384$.

### v3

@@ III-B Scaling Vector Determination
First, we convert $\hat{A}$ and $\hat{B}$ to integer matrices $\hat{A}^{\prime}$ and $\hat{B}^{\prime}$, respectively, as
@@
We consider two strategies: a fast mode and an accurate mode. From $|\hat{a}^{\prime}_{ij}|\leq|\hat{a}_{ij}|\cdot\mu_{i}$ and $|\hat{b}^{\prime}_{ij}|\leq|\hat{b}_{ij}|\cdot\nu_{j}$,
@@
From (6), the row norms of $\hat{A}$ satisfy $\|\hat{a}_{i,:}\|_{2}=\|\hat{a}_{i+m,:}\|_{2}$ for $1\leq i\leq m$, and the column norms of $\hat{B}$ satisfy $\|\hat{b}_{:,j}\|_{2}=\|\hat{b}_{:,j+n}\|_{2}$ for $1\leq j\leq n$. Thus, the scaling values for rows $i$ and $i+m$ of $\hat{A}$ are identical; the same holds for columns $j$ and $j+n$ of $\hat{B}$. Consequently, the scaling vectors reduce to $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$. The elements of $\mu$ and $\nu$ can then be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2]; that is,
@@
For any real number $x\in\mathbb{R}$, $\lfloor\log_{2}|x|\rfloor$ represents the unit in the first place of $x$, and thus it can be obtained directly using the $\mathtt{ilogb}(\cdot)$ function. Following [uchino_ozaki2], the quantity $(\log_{2}(P-1)-1)/2-1$ is precomputed as $\mathcal{P}^{\prime}_{\mathit{fast}}:=\mathrm{single}(\log_{2}(P-1)/2-1.5)$. Furthermore, $\log_{2}\|\hat{a}_{i,:}\|_{2}$ and $\log_{2}\|\hat{b}_{:,j}\|_{2}$ can be upper-bounded by $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{a}_{ih}^{2})$ and $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{b}_{hj}^{2})$, respectively, using floating-point arithmetic in round-up mode. Here, $\mathtt{\_\_log2f}$ denotes the fast built-in base-2 logarithm function provided by the CUDA Math API and the HIP math API. According to NVIDIA’s official CUDA documentation, the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is $2^{-22}$ for $x\in[0.5,2]$, and at most 2 units in the last place (ULPs) otherwise [cuda]. The definition of ULP used in the CUDA documentation follows Muller [muller:inria-00070503]. In addition, the AMD HIP documentation states that the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is 1 ULP for $x\in[10^{-6},10^{6}]$ [hip], although the precise definition of ULP is not specified. Based on these specifications, we assume that the bound $\mathtt{\_\_log2f}(x)\leq(\log_{2}x)/(1-4u)$, where $u$ is the unit roundoff of FP32 arithmetic. If $x$ is not sufficiently close to $1$, it is first scaled to $x^{\prime}:=x/2^{\lfloor\log_{2}|x|\rfloor}$, after which the logarithm is evaluated as $\mathtt{\_\_log2f}(x^{\prime})+\lfloor\log_{2}|x|\rfloor$. Based on the above considerations, for $\delta:=\mathrm{single}(0.5/(1-4u))$ computed in round-down mode, the scaling vectors $\mu$ and $\nu$ are determined as follows:
@@
where $\mathcal{P}^{\prime}_{\mathit{fast}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using the fused multiply-add operation in round-down mode.
In the accurate mode, we first apply diagonal scaling to $A_{R}$, $A_{I}$, $B_{R}$, and $B_{I}$ so that the integer parts of the maximum absolute values in each row of $A_{R}$ and $A_{I}$ and in each column of $B_{R}$ and $B_{I}$ fit within 6 bits. Specifically, we scale them as $\mathrm{diag}(\bar{\mu})A_{R}$, $\mathrm{diag}(\bar{\mu})A_{I}$, $B_{R}\mathrm{diag}(\bar{\nu})$, and $B_{I}\mathrm{diag}(\bar{\nu})$, and then round them up using the ceiling operation to obtain 7-bit or smaller integer matrices $\bar{A}_{R}$, $\bar{A}_{I}$, $\bar{B}_{R}$, and $\bar{B}_{I}$, respectively. These matrices satisfy
@@
where no computing error occurs in $\bar{A}_{R}-\bar{A}_{I}$ and $\bar{B}_{R}-\bar{B}_{I}$ in INT8. Then, the elements of $\mu$ and $\nu$ can be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2]; that is,
@@
where $\mathcal{P}^{\prime}_{\mathit{accu}}:=\mathrm{single}(\log_{2}(P-1)/2-0.5)$ and $\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using the fused multiply-add operation in round-down mode.
@@ III-C Performance Analysis
To analyze the performance characteristics of the proposed methods, we employ a simple yet expressive performance model that captures both computation and memory transfer costs. Let $b$ denote the peak sustained memory bandwidth (in B/s) and $p$ denote the peak throughput of the INT8 matrix-multiplication engine (in operations per second, OPS). For a given operation, the memory-bound execution time is estimated as the required data movement divided by $b$ and the compute-bound time is obtained as the arithmetic workload divided by $p$. Because even memory-bound kernels incur a non-negligible amount of arithmetic overhead, we include a correction term $c$.
@@ III-C1 Fast Mode
Let the input matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$ be double-precision. In step 1, both $A$ and $B$ are loaded twice: once for computing the row-wise maximum absolute values of $\hat{A}$ and the column-wise maximum absolute values of $\hat{B}$, and once for computing the residues of $\hat{A}$ and $\hat{B}$. Afterward, the scaling vectors and the INT8 matrices are stored. Each scaling factor is a power of two, so only its exponent needs to be stored as an INT16. Including the correction term $c$, which accounts for arithmetic overhead even in memory-bound operations, the total execution time for step 1 can be estimated as $((32+3N+c)k+2)(m+n)/b$. In step 2, the INT8 GEMM kernels are executed $3N$ times. The resulting $3N$ INT32 matrices are then loaded and converted into $N$ INT8 complex matrices. Accounting for computational overhead $c$, the execution time for step 2 is estimated as $6Nmnk/p+(14N+c)mn/b$. In step 3, the $N$ INT8 complex matrices are loaded and the CRT accumulation and final reduction are performed. The scaling vectors are then loaded, the inverse scaling is applied, and finally the resulting double-precision complex matrix is stored. With the correction term $c$, the execution time for step 3 is estimated as $(2(m+n)+(2N+16+c)mn)/b$.
Combining the costs derived for steps 1, 2, and 3, the total execution time for the proposed methods can be estimated as the sum of its memory-bound and compute-bound components. Aggregating all terms yields the following expression for the overall execution time:
@@ III-C2 Accurate Mode
and for the single-precision case as
@@
Figs. 2 and 3 show performance model heatmaps for the fast and accurate modes in single and double precision for $6$ and $13$ moduli, respectively. The TFLOPS was computed as
@@
where $t$ is the execution time in seconds. For example, assuming an effective memory bandwidth of 2–4 TB/s and an observed INT8 GEMM kernel throughput of 1500 TFLOPS on the GH200, the model predicts that ZGEMM emulation in the accurate mode would achieve approximately 120 TFLOPS.
Figure 2: Performance model heatmaps for single-precision complex matrix multiplication emulation. The left panel corresponds to the fast mode and the right panel corresponds to the accurate mode. The horizontal axis denotes the achievable memory bandwidth and the vertical axis denotes the achievable INT8 GEMM throughput. The problem size is fixed at $m=n=k=16384$ and the correction term is set to $c=6$, equal to the number of moduli used in the emulation. The color scale indicates the predicted throughput (in TFLOPS) of the proposed emulation.
Figure 3: Performance model heatmaps for double-precision complex matrix multiplication emulation. The left panel corresponds to the fast mode and the right panel corresponds to the accurate mode. The horizontal axis denotes the achievable memory bandwidth and the vertical axis denotes the achievable INT8 GEMM throughput. The problem size is fixed at $m=n=k=16384$ and the correction term is set to $c=13$, equal to the number of moduli used in the emulation. The color scale indicates the predicted throughput (in TFLOPS) of the proposed emulation.
@@ III-D GEMM Emulation Library
In this study, we developed a library that provides a complete emulation environment for GEMM operations based on the Ozaki-II scheme. Details of the library and usage instructions are provided in [GEMMul8]. The library supports both CUDA and HIP backends. The CUDA-based source code is automatically translated into HIP C++ when compiled under the HIP backend. The implementation provides emulation routines for single- and double-precision GEMM, proposed in previous work [uchino_ozaki2], as well as the CGEMM and ZGEMM emulation kernels developed in the present study. The supported data types are float, double, cuComplex, cuDoubleComplex, hipFloatComplex, and hipDoubleComplex.
To enable transparent integration with existing high-performance computing applications, the library can be activated by adding the path of the shared library to the environment variable LD_PRELOAD. Once activated, the following GEMM calls to CUDA or HIP backends are dynamically redirected to the emulation routines, allowing users to evaluate GEMM emulation without modifying their applications:
@@
The implementation consists of INT8 matrix multiplications, fixed-order basic arithmetic operations, and the evaluation of the $\log_{2}(\cdot)$ function in (11), (12), (13), and (14). Rounding errors arise only in the computation of the scaling vectors defined by (11), (12), (13), and (14), and in the accumulation and final reduction stages of the CRT. Except for the logarithm evaluation, all computations are deterministic sequences of basic arithmetic, and thus the emulation produces bitwise-reproducible results as long as the precision of the $\log_{2}(\cdot)$ computation is identical. When the API versions coincide, each backend employs identical implementations of __log2f, ensuring reproducibility across runs.
@@ IV Numerical Results
Here, we compare the following methods:
@@
hipBLAS does not provide hipblasCgemm3m or hipblasZgemm3m; the same limitation applies to rocBLAS, though it is not used in this study. We empirically confirmed that both the performance and numerical accuracy of cublasCgemm3m are essentially identical to those of cublasCgemm; thus, we omit the former’s results from the following comparisons. The cuBLAS implementation of Ozaki-I-based emulation provides two execution modes: Dynamic Mantissa Control and Fixed Mantissa Control. In the Dynamic Mantissa Control mode, the routine determines the number of mantissa bits, or equivalently the number of slices, required for fixed-point emulation such that the numerical accuracy matches or exceeds that of FP64. If the required mantissa length exceeds either an internal limit or a user-specified maximum, the computation is automatically executed in native FP64. In the Fixed Mantissa Control mode, the user specifies the mantissa bit length used for the fixed-point representation. This mode reduces runtime overhead by eliminating dynamic precision selection, but it does not ensure FP64-level accuracy. In this study, the Fixed Mantissa Control mode is used to enable a direct comparison of computational throughput and to quantify the performance gain attributable to fixed-precision emulation.
@@ IV-A Accuracy
where $\mathrm{rand}\in(0,1]$ denotes a uniform random number and $\mathrm{randn}$ denotes a standard normal random number. Both types of random number were generated using either the cuRAND API or the hipRAND library. Following the methodology adopted in prior studies on Ozaki schemes [ozaki2012error, ozaki2013generalization, mukunoki2020, ootomo2024dgemm, uchino2025Performance, ozaki-scheme2, uchino_ozaki2], we introduce the parameter $\phi$ to control the dynamic range of the input matrices. In accordance with these studies, we specify $\phi\in\{0,0.5,1,1.5\}$ for single-precision tests and $\phi\in\{0.5,1,2,4\}$ for double-precision tests. As $\phi$ increases, the maximum magnitude grows exponentially, whereas the median and interquartile range remain comparatively stable. This confirms that $\phi$ effectively widens the dynamic range without substantially affecting the central tendency of the distribution. Such controlled variation is essential for evaluating the performance and numerical stability of Ozaki schemes.
Figs. 4 and 5 show the accuracy of CGEMM and ZGEMM, respectively, observed on the GH200. Similar results were obtained on other GPUs. The maximum relative error in the figures was calculated as
@@
In CGEMM emulation, fast-$N$ with $6\leq N\leq 9$ and accu-$N$ with $6\leq N\leq 8$ provided accuracy comparable to that of the native CGEMM. In ZGEMM emulation, fast-$N$ with $13\leq N\leq 18$ and accu-$N$ with $13\leq N\leq 17$ achieved accuracy comparable to that of the native ZGEMM. As $\phi$ increases, the range of exponents for $A$ and $B$ expands, making it easier to overestimate the upper bound $2\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ in condition (4). DGEMM emulation in [uchino_ozaki2] required at least 14 moduli to achieve native DGEMM-level accuracy, whereas the proposed method attains native ZGEMM-level accuracy with only 13 moduli. This reduction is attributed to the use of the Karatsuba algorithm in the INT8 complex matrix multiplication, which suppresses precision loss more effectively than floating-point arithmetic.
@@
Figure 4: Maximum relative error of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=n=1024$ and $k=16384$.
Figure 5: Maximum relative error of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=n=1024$ and $k=16384$.